a spherical bearing is to have a circumference of 7.2 cm with an error of no more than 0.1 cm. Set up an absolute value inequality and then determine the acceptable range of values for the diameter of the bearing.
Answer by jim_thompson5910(35256)

Let x = diameter of any random spherical bearing


If the circumference is 7.2, then the diameter is d = 7.2/pi = 2.2918

The absolute value inequality we'll use is

|x - t| <= e

where

t = target
e = error

So the target is roughly 2.2918 and the error is 0.1 (or less), so t = 2.2918 and e = 0.1

|x - t| <= e

|x - 2.2918| <= 0.1

-0.1 <= x - 2.2918 <= 0.1

-0.1 + 2.2918 <= x <= 0.1 + 2.2918

2.1918 <= x <= 2.3918

So the acceptable range of values for the diameter is between 2.1918 and 2.3918 (inclusive)